HUBBLE SPACE TELESCOPE DAILY REPORT #817 PERIOD COVERED: 0800 EST 02/04/93 - 0800 EST 02/05/93 Daily Status Report as of 036/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WFC/PC 3172 (Observations: Internal Monitor Proposal) Both the Wide Field Camera and the Wide Field Camera were used to take internal flat and dark calibration observations. No external targets were used. The observations were received and displayed. One of the images was reported to be underexposed. 1.2 Completed WFC 3917 (HST Medium Deep Survey Part 1) The Wide Field Camera was used to take four medium-deep high latitude images taken in parallel with the prime science observations. All four observations were received and displayed in the OSS. The third observation showed a satellite trail across three of the chip images 1.3 Completed PC 4140 (Rapid Internal Monitor Part 2) The Planetary Camera was used to make two internal flat observations to check for new forms of contamination and the return of quantum efficiency hysteresis. Both images were displayed, and reported to look okay. 1.4 Completed WFC 4105 (HST Medium Deep Survey, High Latitude) The Wide Field Camera was used in parallel with the prime science observations to take one medium deep high latitude sky survey image. The observations was received and displayed. It was reported to look okay. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 Ephemeris Uplinked at 036/0152Z. 1.6 Completed GHRS/2 4251 (Physical Conditions and Velocity Structure in Red Giants) The High Resolution Spectrograph was used to make six science spectral observations of the hot red giant binary star EG Andromedae (aka,HD4174) to help fine the mechanism for the star's high loss of mass. The observations was made during a LOS, and still remain on the science tape recorder. They will be analyzed later. 1.7 Completed GHRS/2 3614 (Boron as a Probe of Stellar Structure) The High Resolution Spectrograph was used to study the boron content of star HD27697 to help study the processes in its interior. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 1160 (Absorption Cl. Phys. in Seyfert Galactic Nuclei) 3.2 Complete GHRS/2 4012 (Dark Count Statistics for GHRS Detectors) 3.3 Complete Two Sets of GHRS/2 3614 (Boron as a Probe of Stellar Structure) 3.4 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.5 Complete PC 3230 (PC PSF and Focus Monitoring 2) 3.6 Complete HSP/UV1 4247 (UV Observations of Geminga) 3.7 Complete Two Sets of WFC/PC 3172 (WF/PC SV Observations: Internal Monitor Proposal) 3.8 Complete PC 4347 (WF/PC Calibration: Flood Retention Test) 3.9 Complete PC 4712 (WF/PC Calibration: Measles Monitor) 3.9 Complete GHRS/2 3626 (Ephemeral Deter. of Wind Velocity and Density) 3.10 Complete GHRS/2 3737 (the Onset of Chromospheric Activity) 3.11 Complete FOS/BL 3268 (UV Spectra of QSOs with Z > 3.1 Observations) 3.12 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.13 Complete GHRS/2 3934 (Physical Conditions and Velocity Structures) 3.14 Complete FOS/RD 3840 ( Abundances and Evolution of Carbon Nitrogen and Oxygen) 3.15 Complete WFC 3917 (HST Medium Deep Survey Part 1) 3.16 Complete FGS 2941 (High Speed Astrometry Search for Planet Companions) 3.17 Complete FOC 3790 (Polarization Imaging of Radio Galaxies) 3.18 Complete FOC/96 3801 (QSOs Suitable for Subsequent Observations) 3.19 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.20 Initiate FOC/96 3487 (Cooling Flows in Distant Clusters) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-